kaylablunt0008:

Use the discriminant to describe the roots of each equation. Then select the best description. x2 + 9x + 14 = 0 double root real and rational roots real and irrational roots non-real roots

Nnesha:

\[\rm Given~~ax^2+bx+C=0\]\[\rm Discriminant= b^2-4ac\] IF\[\rm b^2-4ac<0~~imaginary ~roots/no-real~roots\]\[\rm b^2-4ac >0 ~~2~real ~roots\] \[\rm b^2-4ac =0~(1~real~root)\]